1. What is the magnitude of the electric force of attraction between an iron nucleus (q = +26e) and its innermost electron if the distance between them is 1.5 x 10–12 m?
1
Expert's answer
2020-04-20T10:40:51-0400

Use the expression for Coulomb's force:


"F=\frac{1}{4\pi\epsilon_0}\frac{Qq}{r^2}=\\
\space\\
=\frac{1}{4\pi\epsilon_0}\frac{26e\cdot e}{r^2}=\\
\space\\
=\frac{1}{4\pi\epsilon_0}\frac{26\cdot(1.6\cdot10^{-19})^2}{(1.5\cdot10^{-12})^2}=2.7\text{ N}.\\"